# Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal

### v1

@@ Appendix C Performance under Different Token Budgets
Appendix D Generalization to Different Architectures
To evaluate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three baselines: Zero-shot, Original, and SPIRIT. Results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, and LeetCodeDataset benchmarks are shown in Table 12.
Methods HumanEval LiveCodeBench LeetCodeDataset
HE HE+ v1_v3 v4_v5 All
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 68.29 3334 1.86 64.02 3334 1.86 44.12 7162 6.92 25.00 8508 8.90 27.19 8358 8.65
Original 82.32 2978 1.63 76.22 2978 1.63 52.61 6614 6.16 31.34 8202 8.60 26.32 8413 8.85
SPIRIT 76.22 3159 1.74 72.56 3159 1.74 52.61 6280 5.84 30.22 7913 8.45 26.75 8449 8.73
ASAP 80.49 2494 1.30 76.83 2494 1.30 48.86 3605 2.18 32.84 4175 2.69 27.63 3792 2.42
Table 12: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix E Prompts Used in Our Method
Prompt for Direct CoT Generation Given a question, please tell me how to get this answer step by step. Question: ``` {question} ``` Answer: ```python {answer} ``` Only return a detailed (with code) step-by-step solution (containing only “Step-by-Step Solution” and “Final Code”). The detailed step-by-step solution is:

### v2

@@ Appendix F Performance under Different Token Budgets
Budget Zero-shot Original SPIRIT ASAP
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
1K 48.75 963 0.19 59.29 942 0.19 54.66 925 0.18 83.93 693 0.14
2K 83.55 1301 0.27 86.35 1250 0.26 88.55 1118 0.23 90.75 753 0.16
4K 88.65 1553 0.37 90.37 1432 0.34 90.52 1227 0.28 91.28 778 0.18
6K 89.23 1714 0.46 91.05 1513 0.39 91.28 1297 0.33 91.81 790 0.20
Table 15: Results of different methods under different budgets on GSM8K. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Budget Zero-shot Original SPIRIT ASAP
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
1K 19.00 1020 0.19 28.00 1017 0.19 18.40 1012 0.19 36.40 935 0.19
2K 42.20 1804 0.39 52.00 1767 0.39 54.40 1592 0.36 59.80 1347 0.31
4K 60.40 2629 0.70 63.80 2511 0.66 64.20 2144 0.57 70.80 1649 0.43
6K 66.60 3100 0.94 70.60 2843 0.84 69.60 2460 0.74 71.00 1758 0.52
Table 16: Results of different methods under different budgets on MATH500. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
. Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 0.00 2048 0.93 16.67 2048 0.93 6.67 2048 0.93 10.00 1978 0.93 4K 20.00 4003 2.12 20.00 3984 2.11 23.33 3877 2.10 36.67 3415 1.99 6K 33.33 5682 3.58 36.67 5695 3.55 40.00 5216 3.32 36.67 4410 3.03 8K 30.00 7073 5.30 40.00 7093 5.14 46.67 6243 4.47 40.00 5159 4.10 10K 36.67 8352 6.76 46.67 8034 6.43 46.67 7198 5.78 46.67 5552 5.04 12K 40.00 9318 7.84 46.67 8990 7.75 46.67 8363 7.21 46.67 5767 5.88
Table 17: Results of different methods under different budgets on AIME24. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
. Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 6.67 2046 0.95 13.33 2048 0.94 3.33 2044 0.94 10.00 2020 0.95 4K 20.00 3851 2.17 26.67 3834 2.16 16.67 3792 2.15 20.00 3511 2.15 6K 30.00 5369 3.62 33.33 5452 3.63 36.67 5360 3.63 30.00 4484 3.14 8K 36.67 6848 5.30 36.67 6798 5.32 36.67 6611 5.16 33.33 5002 4.10 10K 40.00 8145 6.67 43.33 8026 6.75 43.33 7817 6.57 36.67 5434 5.10 12K 36.67 9442 8.25 40.00 9461 8.32 46.67 8598 7.78 36.67 5720 6.06
Table 18: Results of different methods under different budgets on AIME25. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix G Training Efficiency
To quantify the training efficiency gains, we present results of the CodeForces-CoTs dataset in Table 5 and results of the OpenR1-Math dataset in Table 19. We report two key metrics: the average number of tokens per sample and the average training time measured in seconds per step.
Methods Tokens Time
Original 5807 47.82
Selective Context 3149 (-45.8%) 25.85 (-45.9%)
LLMLingua-2 3478 (-40.1%) 28.75 (-39.9%)
TokenSkip 4728 (-18.6%) 39.20 (-18.0%)
SPIRIT 2858 (-50.8%) 23.67 (-50.5%)
ASAP 1834 (-68.4%) 15.36 (-67.9%)
Table 19: Training efficiency comparison on OpenR1-Math dataset. We report the average number of tokens per sample and training time measured in seconds per step. Percentages indicate the reduction relative to the Original baseline.